• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Adaq

    Log entry time 21:53:54 on June 8,2002

    Entry number 85565

    keyword=Left Arm Motion

    Spectrometer motions fixed: Bogie Left VME crate had to be reset (green buttons) in middle room